# REFINER: Reasoning Feedback on Intermediate Representations

### v1

@@ Abstract
Language models (LMs) have recently shown remarkable performance on reasoning tasks by explicitly generating intermediate inferences, e.g., chain-of-thought prompting. However, these intermediate inference steps may be inappropriate deductions from the initial context and lead to incorrect final predictions. Here we introduce REFINER, a framework for finetuning LMs to explicitly generate intermediate reasoning steps while interacting with a critic model that provides automated feedback on the reasoning. Specifically, the critic provides structured feedback that the reasoning LM uses to iteratively improve its intermediate arguments. Empirical evaluations of REFINER on three diverse reasoning tasks show significant improvements over baseline LMs of comparable scale. Furthermore, when using GPT3.5 as the reasoner, the trained critic significantly improves reasoning without finetuning the reasoner. Finally, our critic model is trained without expensive human-in-the-loop data but can be substituted with humans at inference time.
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language processing (NLP) tasks (Brown et al., 2020). Recent work has shown that explicitly generating intermediate steps during reasoning tasks significantly improves a model’s performance and interpretability (Shwartz et al., 2020; Paul and Frank, 2021; Marasovic et al., 2022; Lampinen et al., 2022; Wei et al., 2022). Producing such intermediate representations provides insight into the model’s predictions and allows humans to inspect the model’s reasoning process. However, these intermediate representations [^1] In a reasoning task, the intermediate representations can be viewed as inference rules, explanations or reasoning steps. can be unreliable (Ye and Durrett, 2022) and result in poor performance on downstream reasoning tasks. Most importantly, it is unclear how to meaningfully refine the intermediate representations to further improve the final performance.
@@
The standard practice for correcting reasoning errors is to annotate new data and either retrain or finetune the model (Feng et al., 2021; Hedderich et al., 2021). However, fixing such errors by finetuning with more data is not only data- and resource-intensive but can also be insufficient to generalize well in complex reasoning tasks (Ward et al., 2022). Other works have explored improving models using feedback by providing a scalar score (Ziegler et al., 2019; Martin et al., 2022) or directly revealing the correct missing answer (Mehta and Goldwasser, 2019; Elgohary et al., 2021; Tandon et al., 2022). However, in natural language reasoning tasks, defining a scalar value that captures different fine-grained reasoning error types (e.g., semantic consistency, logical, etc.) remains an open challenge (Golovneva et al., 2023).
@@
To provide fine-grained feedback about reasoning errors, we develop a scheme to independently train the critic model on automatically constructed feedback data. More specifically, we create pairs of incorrect intermediate representations and structured [^2] Note that we transform the structured feedback into semi-structured textual feedback using templates. feedback on fine-grained reasoning errors. Then, we use this data to train the critic to provide fine-grained feedback on erroneous intermediate reasoning steps. Finally, the critic interacts with the generator LM, offering feedback both during the training of the generator and during inference.
@@
Contributions.
(i) We propose REFINER, a framework that refines LMs reasoning capabilities through feedback. To the best of our knowledge, our work is the first to investigate how interacting with fine-grained reasoning feedback on intermediate reasoning steps impacts the performance of LMs on reasoning tasks.
(ii) We evaluate REFINER on three natural language reasoning tasks: math word problems (Koncel-Kedziorski et al., 2016; Patel et al., 2021), synthetic natural language reasoning (Liang et al., 2022), and moral action generation (Emelin et al., 2021). REFINER demonstrates significant performance gains across different LM architectures with different scales. Across different reasoning tasks, REFINER outperforms comparably-sized strong fine-tuned LM baselines (by +13.1, +3.2, +15 pts., respectively).
(iii) We empirically demonstrate that for math word problem and synthetic natural language reasoning, our trained critic models alone are beneficial for improving intermediate representations as they help GPT3.5 significantly increase its performance in a few-shot setting (by +3, +9.2 pts., respectively). We also demonstrate that providing structured feedback on fine-grained errors can benefit more than scalar value feedback for moral action generation and math word problem tasks. Our critic model acts as a ‘reasoning refinement tool’ for LLMs.
(iv) Our analyses show that (a) improving the intermediate representation generation improves the performance on the reasoning tasks, and (b) training a generator with an imperfect (noisy) critic is still beneficial.
Our code is made publicly available  [^3] https://github.com/debjitpaul/refiner
@@ 2 Related Work
Intermediate Representations.
While state-of-the-art LMs achieve incredible performances in a wide range of tasks, they have difficulty with many reasoning tasks Wang et al. (2022), especially ones with multiple constraints or sub-problems, similar keywords to more common tasks, or requiring specialized knowledge Austin et al. (2021) – such as mathematical problem solving Ling et al. (2017); Andor et al. (2019); Ran et al. (2019); Geva et al. (2020); Piękos et al. (2021); Cobbe et al. (2021); Kim et al. (2022).
@@
Natural Language Feedback.
Recent work has explored giving models richer and more complex feedback through the use of natural language Ziegler et al. (2019); Nguyen et al. (2021); Scheurer et al. (2022), used for aligning LLMs’ output with users’ preferences Christiano et al. (2017); Ziegler et al. (2019); Saunders et al. (2022); Scheurer et al. (2022); Bai et al. (2022), or to directly improve the model’s performance in its current task Weston (2016); Rupprecht et al. (2018); Elgohary et al. (2020); Austin et al. (2021); Madaan et al. (2023).
This training is usually dependent on human-created feedback, generated in large quantities Bai et al. (2022), which takes up considerable resources. Though an external feedback provider can guide models to correct answers and reasoning Austin et al. (2021), demonstrably better than they can themselves Saunders et al. (2022), feedback has rarely been used in this way – and automated critics for reasoning tasks have proved to be difficult Scheurer et al. (2022); Wang et al. (2022); Huang et al. (2022).
@@ 3 REFINER
Problem Formulation. In this paper, we view natural language reasoning (NLR) as an autoregressive generation task where, given input context $x$, a model needs to generate $y$, such that $y$ satisfies the constraints of the task. Usually, to generate correct or plausible $y$, the model needs to make the correct inference $z$ as intermediate steps. [^4] We use “inference steps/representations” and “hypothesis” interchangeably. We decompose NLR tasks as follows:
$p(y|x)=p(y|x,z)p(z|x)$ (1)
In practice, one can compute each conditional using an LM that includes its conditioning variables as a part of its input.
Before continuing with the model description, we illustrate three NLR tasks where we conduct our study and their respective intermediate representation $z$. These three tasks broadly cover two types: (i) logical reasoning and (ii) normative reasoning. They are exemplified in Figure 2 and detailed below.
Synthetic natural language reasoning (sNLR), where given a reasoning scenario $x$ consisting of $5$ synthetic rules and a fact, the model needs to deduce a conclusion $y$. This task requires the model to perform deductive reasoning and generate the conclusion $z$ using closed-world rules and facts.
@@
We propose to solve these tasks by forcing the model to generate intermediate hypotheses ($z$) and improving them via structured feedback. We introduce an interactive framework named REFINER, made of two separate models: (a) a critic model (§3.1) trained to provide structured feedback on intermediate reasoning steps and (b) a generator model trained to solve the reasoning task by first generating intermediate reasoning steps (§3.2). The core idea of REFINER is to exploit the interaction between the generator model and the critic model, where the generator’s intermediate reasoning steps are improved via structured feedback from the critic. An overview of the framework is depicted in Fig. 3.
Figure 2: An overview of the three tasks tackled in this paper, with examples of both valid and invalid intermediate reasoning steps, as well as their corresponding fine-grained error types. Notice the Missing Steps error type, in the second task, actually encompasses two error types: reasoning misalignment, derived from not considering the or operation, and lack of implicit knowledge, where implicit knowledge is needed to match the existing rules.
@@
Incorrect Numbers The position number
in equation-number
is incorrect.
@@
equation-number
is incorrect.
@@
Logically Invalid The X operator
makes inference rule
number invalid.
sNLR Missing Link Missing link between
the fact and the rules.
Missing Implicit The implicit knowledge
Knowledge Step is missing.
@@
MS Semantic Misalignment Semantically misaligned:
“ text snippet”
@@
3.1 CRITIC Model
The role of the critic is to provide feedback on the intermediate hypotheses produced by the generator model. One way to evaluate the quality of the hypothesis and produce feedback on the hypothesis $z$, would be to compare it against a gold hypothesis $z^{*}$. Previous works employed automatic metrics like BLEU, ROUGE, etc., as value functions (Wu et al., 2018; Ramamurthy et al., 2022). However, these scalar value functions are not suitable for natural language reasoning tasks because: (i) it is unclear how to define a scalar value function that can encapsulate fine-grained reasoning errors (Golovneva et al., 2023) and (ii) during inference, these functions require access to the gold hypothesis (which is unavailable in practice). Therefore, we train a critic model and endow it with the ability to evaluate the hypothesis in a fine-grained manner and provide structured feedback.
Feedback Data Generation.
To train the critic, we have to create example pairs of implausible hypotheses and their corresponding feedback with fine-grained reasoning errors. Inspired by Golovneva et al. (2023) and Talmor et al. (2020), we first define fine-grained reasoning error types for each reasoning task (see Table 1). For MWP, an equation can be incorrect due to: (i) the operands or operators in the equations being incorrect and/or (ii) one or more operators missing. For sNLR, an inference rule can be incorrect because it is (i) logically invalid and/or (ii) missing reasoning rules (failing to connect the correct facts with correct rules or missing implicit knowledge). For MS, a moral norm can be incorrect due to (i) contradiction and/or (ii) semantic misalignment. In Figure 2, we provide a few examples of fine-grained error types for each reasoning task.
@@ 3.1 CRITIC Model
For MWP and sNLR problems, the underlying reasoning requires symbolic systems with closed-world rules. Hence, we consider a simple rule-based method to automatically generate the pairs of errors and their corresponding structured feedback by considering the error types and position of the errors (see Fig. 2).
In the moral norm generation task, we consider two kinds of fine-grained errors: logical contradiction and semantic misalignment (incoherent, uninformative). Moral norms are people’s subjective judgments about the character and actions mentioned in the context. Each moral norm is a combination of two components (implicit structure): a moral judgment [You shouldn’t] and an action [criticize your family’s religion]. Firstly, to create logical contradictions, we use the concept of deontic logic from Kiehne et al. (2022) and derive new norms contrary to those of Moral Stories. Hence, we replace the correct moral judgments in the plausible hypothesis with inverse judgments. For example, replacing [You shouldn’t] to [It’s good] from the plausible hypothesis as depicted in Fig. 2. To scale such inverse norms (implausible hypothesis), we paraphrase them by substituting the adjectives with synonyms from WordNet (see Fig. 2). Secondly, to create semantic misalignments, we should collect implausible hypotheses that are either misaligned with the plausible hypothesis or incomplete in nature. To create them, we replace the correct action (verb phrase) from the plausible hypothesis with random verb phrases selected from the context of the plausible hypothesis.
We also replace the correct judgment with random judgments to scale the number of implausible hypotheses per example. Finally, as feedback $f$, we provide <<error type, hint>>. For non-monotonic reasoning tasks [^5] A non-monotonic reasoning deals with incomplete facts, whereby reasoners draw tentative conclusions, which may be retracted based on further evidence. like norm and action generation, the critic should be able to provide hints that align the generator model’s objective to the reasoning task. Hence, as a hint, we provide verb phrases from the norms. Since the critic provides textual feedback to the generator, we convert the structured feedback into natural language feedback  [^6] Further details about feedback are provided in Appendix.. Formally, we create a data pool $D=\{x,z,z^{\prime},f\}$ to train a critic model.
Training the critic model.
We train a supervised critic model ($\pi_{\beta}$) with the context ($x$) and (plausible or implausible) hypothesis ($z$ or $z^{\prime}$) as input and the textual feedback as output. We update the critic using the following cross-entropy loss:
$L(\beta)=-\log p_{\beta}(f(u)|x,u)$ (2)
where $u\in z,z^{\prime}$. The trained critic is only used during inference. The oracle critic is used while training the generator model.
Figure 3: Overview of REFINER interaction loop. In each iteration, the generator generates multiple hypotheses. The critic randomly selects one hypothesis and provides feedback based on reasoning errors.
@@ 3.2 GENERATOR Model
Warm-up.
Given a context $x$ the generator model ($\pi_{\theta}$) is trained to generate plausible hypotheses. The warm-up phase is critical to ensure that, when the critic comes in the loop, the generator does not produce random answers likely to be very bad, given the size of the output space. Hence, we use a relatively small supervised dataset (10% training data) to fine-tune the model on the NLR task of interest. After the warm-up phase, we use the additional feedback $f$ from the critic model and learn $\pi_{\theta}(z|x,z^{\prime},f)$.
Exploration.
At each iteration ($t$), the generator model generates multiple hypotheses ($z^{k}$) using nucleus sampling. The critic model randomly selects one hypothesis and provides feedback on that hypothesis. The exploration step aims at increasing the output variance such that the generator receives a wide range of feedback during training.
Learning.
We update the generator model using the following cross-entropy loss:
$L(\theta)=-\sum_{t=1}^{T}\log p_{\theta}(z_{t}|x,z_{t}^{\prime},f_{t}(z^{\prime}))$ (3)
where $T$ = total number of iterations. Since the feedback contains the error types and hints, which are (latent) fine-grained and logical, it should allow the model to learn and update its generation by addressing the reasoning errors mentioned in the feedback.
Inference.
We use the trained critic (see Eq. 2) along with the trained generator to generate a trajectory $z_{0},z_{1},...,z_{T}$ and stop when either $f(z_{t})$ is generated by the generator or “No hint” is generated by the critic. We also experimented with chain of thought prompting, where the generator generates a trajectory $z_{0}y_{0},z_{1}y_{1},...,z_{T}y_{T}$ and stops when the critic generates “No hint”.

### v2

@@ Abstract
Language models (LMs) have recently shown remarkable performance on reasoning tasks by explicitly generating intermediate inferences, e.g., chain-of-thought prompting. However, these intermediate inference steps may be inappropriate deductions from the initial context and lead to incorrect final predictions. Here we introduce REFINER, a framework for finetuning LMs to explicitly generate intermediate reasoning steps while interacting with a critic model that provides automated feedback on the reasoning. Specifically, the critic provides structured feedback that the reasoning LM uses to iteratively improve its intermediate arguments. Empirical evaluations of REFINER on three diverse reasoning tasks show significant improvements over baseline LMs of comparable scale. Furthermore, when using GPT-3.5 or ChatGPT as the reasoner, the trained critic significantly improves reasoning without finetuning the reasoner. Finally, our critic model is trained without expensive human-in-the-loop data but can be substituted with humans at inference time.
REFINER: Reasoning Feedback on Intermediate Representations
Debjit Paul, Mete Ismayilzada, Maxime Peyrard, Beatriz Borges, Antoine Bosselut, Robert West, Boi Faltings EPFL Université Grenoble Alpes, CNRS, Grenoble INP, LIG {firstname.lastname}@epfl.ch
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language processing (NLP) tasks (Brown et al., 2020). Recent work has shown that explicitly generating intermediate steps during reasoning tasks significantly improves a model’s performance and interpretability (Shwartz et al., 2020; Paul and Frank, 2021; Marasovic et al., 2022; Lampinen et al., 2022; Wei et al., 2022). Producing such intermediate representations provides insight into the model’s predictions and allows humans to inspect the model’s reasoning process. However, these intermediate representations [^1] In a reasoning task, the intermediate representations can be viewed as inference rules, explanations or reasoning steps.
* Work done at EPFL
can be unreliable (Ye and Durrett, 2022) and result in poor performance on downstream reasoning tasks. Most importantly, it is unclear how to meaningfully refine the intermediate representations to further improve the final performance.
@@
The standard practice for correcting reasoning errors is to annotate new data and either retrain or finetune the model (Feng et al., 2021; Hedderich et al., 2021). However, fixing such errors by finetuning with more data is not only data- and resource-intensive but can also be insufficient to generalize well in complex reasoning tasks (Ward et al., 2022). Other works have explored improving models using feedback by providing a scalar reward (Ziegler et al., 2019; Martin et al., 2022) or directly revealing the correct missing answer (Mehta and Goldwasser, 2019; Elgohary et al., 2021; Tandon et al., 2022). However, in natural language reasoning tasks, defining a reward that captures different fine-grained reasoning error types (e.g., semantic consistency, logical, etc.) remains an open challenge (Golovneva et al., 2023). Additionally, such a reward provides a relatively sparse training signal.
@@
To provide fine-grained feedback about reasoning errors, we develop a scheme to independently train the critic model on automatically constructed feedback data. More specifically, we create pairs of incorrect intermediate representations and structured [^2] Note that we transform the structured feedback into semi-structured textual feedback using templates. feedback on their fine-grained reasoning errors. Then, we use this data to train the critic to provide fine-grained feedback on erroneous intermediate reasoning steps. Finally, the critic interacts with the generator LM, offering feedback both during the training of the generator and during inference.
@@
Contributions. (i) We propose REFINER, a framework that refines LMs reasoning capabilities through feedback. Our work investigates how interacting with fine-grained reasoning feedback on intermediate reasoning steps impacts the performance of LMs on reasoning tasks. We evaluate REFINER on three natural language reasoning tasks: math word problems, synthetic natural language reasoning, and moral action generation. REFINER demonstrates significant performance gains across different LM architectures with different scales. Across different reasoning tasks, REFINER outperforms comparably-sized strong fine-tuned LM baselines (by +13.1, +3.2, +15 pts., respectively). (ii) We empirically demonstrate that for math word problems and synthetic natural language reasoning, our trained critic models alone are beneficial for improving intermediate representations as they help GPT-$3.5$ significantly increase its performance in a few-shot setting (by +3.5, +6.8 pts., respectively). We also demonstrate that providing structured feedback on fine-grained errors can benefit more than scalar value feedback for moral action generation and math word problem tasks. Our critic model acts as a ‘reasoning refinement tool’ for LLMs. (iii) We show that REFINER can substantially outperform other refinement methods that use feedback from large LMs, such as self-refine. (iv) Our analyses illustrate that (a) improving the intermediate representation generation improves the performance on the reasoning tasks, and (b) training a generator with an imperfect (noisy) critic is still beneficial. Our code is made publicly available  [^3] https://github.com/debjitpaul/refiner.
@@ 2 Related Work
Intermediate Representations. While state-of-the-art LMs achieve incredible performances in a wide range of tasks, they have difficulty with many reasoning tasks Wang et al. (2022), especially ones with multiple constraints or sub-problems or requiring specialized knowledge Austin et al. (2021) – such as mathematical problem solving Ling et al. (2017); Andor et al. (2019); Ran et al. (2019); Geva et al. (2020); Piękos et al. (2021); Cobbe et al. (2021a); Kim et al. (2022).
@@
Natural Language Feedback. Recent work has explored giving models richer and more complex feedback through the use of natural language Ziegler et al. (2019); Nguyen et al. (2021); Scheurer et al. (2022), used for aligning LLMs’ output with users’ preferences Christiano et al. (2017); Ziegler et al. (2019); Saunders et al. (2022); Scheurer et al. (2022); Bai et al. (2022), or to directly improve the model’s performance in its current task Weston (2016); Rupprecht et al. (2018); Elgohary et al. (2020); Austin et al. (2021); Madaan et al. (2023). This training depends on human-created feedback, generated in large quantities Bai et al. (2022), which takes up considerable resources. Though an external feedback provider can guide models to correct answers and reasoning Austin et al. (2021), demonstrably better than they can themselves Saunders et al. (2022), feedback has rarely been used in this way – and automated critics for reasoning tasks have proved to be difficult Scheurer et al. (2022); Wang et al. (2022); Huang et al. (2022).
@@ 3 REFINER
Problem Formulation. In this paper, we view natural language reasoning (NLR) as an autoregressive generation task where, given input context $x$, a model needs to generate $y$, such that $y$ satisfies the constraints of the task. Usually, to generate correct or plausible $y$, the model needs to make the correct inference $z$ as intermediate steps. [^4] We use “inference steps/representations” and “hypothesis” interchangeably. We decompose NLR tasks as follows: $p(y|x)=p(y|x,z)p(z|x)$. In practice, one can compute each conditional using an LM that includes its conditioning variables as a part of its input.
Before continuing with the model description, we describe three NLR tasks where we conduct our study and their respective intermediate representation $z$. We deliberately chose these three tasks since they broadly cover two types of reasoning: (i) logical reasoning and (ii) normative reasoning. They are exemplified in Appx  Fig. 6 and detailed below.
@@
Synthetic natural language reasoning (sNLR), where given a reasoning scenario $x$ consisting of $5$ synthetic rules and a fact, the model needs to deduce a conclusion $y$. This task requires the model to perform deductive reasoning and generate intermediate steps $z$ and the conclusion $y$ using closed-world rules and facts.
@@
We propose to solve these tasks by forcing the model to generate intermediate hypotheses ($z$) and improving them via structured feedback. We introduce an interactive framework, REFINER, made of two separate models: (a) a critic model (§3.1) trained to provide structured feedback on intermediate reasoning steps and (b) a generator model trained to solve the reasoning task by first generating intermediate reasoning steps (§3.2). The core idea of REFINER is to exploit the interaction between the generator model and the critic model, where the generator’s intermediate reasoning steps are improved via structured feedback from the critic.
@@
3.1 CRITIC Model
The role of the critic is to provide feedback on the intermediate hypotheses produced by the generator model. One way to evaluate the quality of the hypothesis and produce feedback on the hypothesis $z$, would be to compare it against a gold hypothesis $z^{*}$. Previous works employed automatic metrics like BLEU, ROUGE, etc., as value functions (Wu et al., 2018; Ramamurthy et al., 2022). However, these scalar value functions are not suitable for natural language reasoning tasks because (i) it is unclear how to define a scalar value function that can encapsulate fine-grained reasoning errors (Golovneva et al., 2023) and (ii) during inference, these functions require access to the gold hypothesis (which is unavailable in practice). Therefore, we train a critic model and endow it with the ability to evaluate the hypothesis in a fine-grained manner and provide structured feedback.
Feedback Data Generation. To train the critic, we have to create example pairs of implausible hypotheses and their corresponding feedback with fine-grained reasoning errors. Inspired by Golovneva et al. (2023) and Talmor et al. (2020), we first define fine-grained reasoning error types for each reasoning task (see Table 1). For MWP, an equation can be incorrect due to: (i) the operands or operators in the equations being incorrect and/or (ii) one or more operators missing. For sNLR, an inference rule can be incorrect because it is (i) logically invalid and/or (ii) missing reasoning rules (failing to connect the correct facts with correct rules or missing implicit knowledge). For MS, a moral norm can be incorrect due to (i) contradiction and/or (ii) semantic misalignment.
@@ 3.1 CRITIC Model
Incorrect Numbers The position number in
equation-number is incorrect.
@@
equation-number is incorrect.
@@
Logically Invalid The X operator makes inference
rule number invalid.
sNLR Missing Link Missing link between the fact the rules.
Missing Implicit The implicit knowledge is
Knowledge Step missing.
@@
MS Semantic Misalignment Semantically misaligned: “ text snippet”
@@
Based on these error types, we propose two strategies to create the feedback data: (i) Rule-based perturbation strategy: we perturb the plausible hypotheses ($z$) in the training data and collect a pool of data $D$ ($x$: input, $z$: plausible hypothesis, $z^{\prime}$: implausible hypothesis). We perturb by omitting, replacing or adding some tokens or some rules from the plausible hypothesis to create an implausible hypothesis automatically (details in Appendix F.1). (ii) Synthetic Generation strategy: we prompted OpenAI’s GPT-$3.5$ to generate implausible hypotheses based on the error types automatically. We used a few-shot setting where we varied the instruction, the number of demonstrations, and the formatting of the demonstrations (details in Appendix F.2).
@@
Figure 2: Overview of REFINER interaction loop. Left side: Training the critic model. Right side: In each iteration, the generator generates multiple hypotheses. The critic randomly selects one hypothesis and provides feedback based on reasoning errors.
We also replace the correct judgment with random judgments to scale the number of implausible hypotheses per example. Finally, as feedback $f$, we provide <<error type, hint>>. For non-monotonic reasoning tasks like norm and action generation, the critic should be able to provide hints that align the generator model’s objective to the reasoning task. Hence, as a hint, we provide verb phrases from the norms. Since the critic provides textual feedback to the generator, we convert the structured feedback into natural language feedback  [^5] Further details about feedback are provided in Appx.F.. Formally, we create a data pool $D=\{x,z,z^{\prime},f\}$ to train a critic model.
Training the critic model. We train a supervised critic model ($\pi_{\beta}$) with the context ($x$) and (plausible or implausible) hypothesis ($z$ or $z^{\prime}$) as input and the textual feedback as output. We update the critic with the cross-entropy loss: $L(\beta)=-\log p_{\beta}(f(u)|x,u)$ where $u\in z,z^{\prime}$. The trained critic is only used during inference. The oracle critic is used while training the generator.
@@ 3.2 GENERATOR Model
Warm-up. Given a context $x$ the generator model ($\pi_{\theta}$) is trained to generate plausible hypotheses. The warm-up phase is critical to ensure that, when the critic comes in the loop, the generator does not produce random answers likely to be bad, given the size of the output space. As such, we use a small supervised dataset (10% training data) to fine-tune the model on the NLR task of interest. After the warm-up phase, we use the additional feedback $f$ from the critic model and learn $\pi_{\theta}(z|x,z^{\prime},f)$.
Exploration. At each iteration ($t$), the generator model generates multiple hypotheses ($z^{k}$) using nucleus sampling. The critic model randomly selects one hypothesis and provides feedback on that hypothesis. The exploration step aims at increasing the output variance such that the generator receives a wide range of feedback during training.
Learning. We update the generator model using the following cross-entropy loss: $L(\theta)=-\sum_{t=1}^{T}\log p_{\theta}(z_{t}|x,z_{t}^{\prime},f_{t}(z^{\prime}))$ where $T$ = total number of iterations. Since the feedback contains the error types and hints, which are (latent) fine-grained and logical, it should allow the model to learn and update its generation by addressing the reasoning errors mentioned in the feedback.
Inference. We use the trained critic along with the trained generator to generate a trajectory $z_{0},z_{1},...,z_{T}$ and stop when either $f(z_{t})$ is generated by the generator or “No hint” is generated by the critic. We also experimented with chain of thought prompting, where the generator generates a trajectory $z_{0}y_{0},z_{1}y_{1},...,z_{T}y_{T}$ and stops when the critic generates “No hint”.